# Simulating and investigating various dynamic aspects of $\rm{H}_{2}\rm{O}$-related hydrogen bond model

### v2

@@ IV.1 Comparison between unitary and dissipative evolutions
Dissipation intensities are taken into consideration to calculate the dissipative evolution of an open system with thermal phonon entering and exiting. The QME is numerically calculated using Eq. (22). In addition to initial state $|0\rangle|0\rangle|1\rangle$, the initial state $|1\rangle|1\rangle|0\rangle$ — critical state, is also considered. Critical state means the critical point of hydrogen bond breaking is reached. Comparison between these two initial states is carried out in Fig. 5. As shown in both panels, the periodic oscillations disappear, replaced by dissipations. The time of hydrogen bond formation is about $1\rm{ps}$ ($10^{-12}s$) before the probability reaches a stable level. In panel (a), $\rho_{|-1\rangle|0\rangle|0\rangle}>\rho_{|2\rangle|1\rangle|0\rangle}$, thus the system tends to form a stable hydrogen bond when the initial state is $|0\rangle|0\rangle|1\rangle$. On the contrary, in panel (b), $\rho_{|-1\rangle|0\rangle|0\rangle}<\rho_{|2\rangle|1\rangle|0\rangle}$, the system tends to break the hydrogen bond when the initial state is $|1\rangle|1\rangle|0\rangle$. This is because the state $|0\rangle|0\rangle|1\rangle$ is stretched and ground, and the proton of donor at this time is at lower energy lever. Thus, it is more inclined to form a stable hydrogen bond. The state $|1\rangle|1\rangle|0\rangle$ means that the critical point of hydrogen bond formation and breaking is obtained, and the hydrogen bond breaks more easily to release two free water molecules.
IV.2 The effect of interactions on evolution
@@
Result of Fig. 6 is shown as heat maps, where the color dots represent the probability of stable hydrogen bond formation at $t\to\infty$. Here $\infty$ is obtained, when the system tends to stabilize and probabilities of states approach constant values. In the next subsections, results of Figs. 7, 8 and 11 are also represented as heat maps at $t\to\infty$.
@@
The effect of interaction strengths $g_{dist}$ and $g_{prot}$ on the evolution and the hydrogen bond formation is obtained in Fig. 6, where the values of $g_{dist}$, $g_{prot}$ range from $0.5g$ to $2g$. All dissipation intensities are fixed: $\gamma_{isol}^{out}=\gamma_{bond}^{out}=\gamma_{phn}^{out}=\gamma$. Under different initial states, the dissipative evolution will eventually reach probabilistic stability after long-term evolution.
Figure 7: (online colar) Effect of different dissipations intensities $\gamma_{isol}^{out}$, $\gamma_{bond}^{out}$ and $\gamma_{phn}^{out}$ on the hydrogen bond formation. $g_{prot}=g_{dist}=g$. The first row of panels is the result of the initial state $|0\rangle|0\rangle|1\rangle$. The second row of panels is the result of the initial state $|1\rangle|1\rangle|0\rangle$. The values of $\gamma_{isol}^{out}$, $\gamma_{bond}^{out}$ range from $0.5\gamma$ to $2\gamma$ and the values of $\gamma_{phn}^{out}$ from the first column to the fourth column correspond to $\frac{1}{2}\gamma,\ \gamma,\ \frac{3}{2}\gamma$ and $2\gamma$, respectively. The definitions of regions II, III and IV are all defined in Fig. 6.
@@
In the first row of Fig. 6, when the initial state is $|0\rangle|0\rangle|1\rangle$, result is found that the larger interaction strengths $g_{dist}$ and $g_{prot}$, the lower the probability of stable hydrogen bond formation. However, in the second row, when initial state is $|1\rangle|1\rangle|0\rangle$, the opposite result is found. In summary, in the case of $|0\rangle|0\rangle|1\rangle$, the interaction strengths have the negative effect on the formation of stable hydrogen bond — hindering the hydrogen bond formation; in the case of $|1\rangle|1\rangle|0\rangle$, the interaction strengths have the positive effect on the formation of stable hydrogen bond — promoting the hydrogen bond formation. Besides, when the initial state is $|0\rangle|0\rangle|1\rangle$, the probabilities on the heat maps are all on the range of [0.5, 1]0.51[0.5,\ 1][ 0.5 , 1 ]; when the initial state is $|1\rangle|1\rangle|0\rangle$, the probabilities on the heat maps are all on the range of [0, 0.5)00.5[0,\ 0.5)[ 0 , 0.5 ). This is because according to the results in Fig. 5 (a), when the initial state is $|0\rangle|0\rangle|1\rangle$, the probability of the $|\rm{O}\cdots\rm{H}\rangle$ is greater than $0.8$, and in the first row of Fig. 6 the interaction strengths have slight effect on the evolution, so the probabilities represented by the points on the first row of the heat maps are almost all around $0.8$. Similarly, according to the Fig. 5 (b), when the initial state is $|1\rangle|1\rangle|0\rangle$, the probability of the $|\rm{O}\cdots\rm{H}\rangle$ is less than $0.2$, and in the second row of Fig. 6 the interaction strengths have also slight effect on the evolution, so the probabilities represented by the points on the second row of the heat maps are almost all around $0.2$.
@@
In this section, the effects of the inflows intensities $\gamma_{phn}^{in}$, $\gamma_{bond}^{in}$ and $\gamma_{isol}^{in}$ are investigated. The first column corresponds to the case that inflows are prohibited: $\mu_{isol}=\mu_{bond}=\mu_{phn}=0$, that is to say, $\gamma_{isol}^{in}=\gamma_{bond}^{in}=\gamma_{phn}^{in}=0$. The second column corresponds to the case that only phonon inflow is permitted: $\mu_{isol}=\mu_{bond}=0,\ \mu_{phn}=0.01$. The third column corresponds to the case that all inflows are permitted: $\mu_{isol}=\mu_{bond}=\mu_{phn}=0.01$. Comparing the second and the third columns with the first column, result is found that the addition of weak inflows actually promotes the formation of hydrogen bond, regardless of whether the initial state is $|0\rangle|0\rangle|1\rangle$ or $|1\rangle|1\rangle|0\rangle$. For easy observation, we use the dividing lines. In the first row from panel (a) to panel (c), it is easy to find that the solid dividing line representing the probability of $0.9$ is shifted to the upper right. At the same time, the area of region III of probability with [0.5, 0.9)0.50.9[0.5,\ 0.9)[ 0.5 , 0.9 ) decreases and the area of region IV of probability with [0.9, 1]0.91[0.9,\ 1][ 0.9 , 1 ] increases. This means the weak inflows promotes the hydrogen bond formation when initial state is $|0\rangle|0\rangle|1\rangle$. Similarly, in the second row from panel (d) to panel (f), the solid dividing line representing the probability of $0.1$ is shifted to the lower left. At the same time, the area of region I of probability with [0, 0.1)00.1[0,\ 0.1)[ 0 , 0.1 ) decreases and the area of region II of probability with [0.1, 0.5]0.10.5[0.1,\ 0.5][ 0.1 , 0.5 ] increases. This means the weak inflows also promotes the hydrogen bond formation when initial state is $|1\rangle|1\rangle|0\rangle$.
@@
In addition to the above results, the effect of different interactions strengths on the time required for the system to reach stable is also obtained: the smaller interactions strengths, the more time required to reach stable.
IV.3 The effects of dissipations on evolution
@@
Figure 8: (online colar) Effect of different inflows intensities $\mu_{isol}$, $\mu_{bond}$ and $\mu_{phn}$ on the hydrogen bond formation. $g_{prot}=g_{dist}=g$, $\gamma_{isol}^{out}=\gamma_{bond}^{out}=\gamma_{phn}^{out}=\gamma$. The first row of panels is the result of the initial state $|0\rangle|0\rangle|1\rangle$. The second row of panels is the result of the initial state $|1\rangle|1\rangle|0\rangle$. The values of $\mu_{isol}$, $\mu_{bond}$ range from 00 to $0.9$ and the values of $\mu_{phn}$ from the first column to the fourth column correspond to 0, 0.3, 0.600.30.60,\ 0.3,\ 0.60 , 0.3 , 0.6 and $0.9$, respectively. The definitions of regions I, II, III and IV are all shown in Fig. 6.
In the first row of Fig. 7, when the initial state is $|0\rangle|0\rangle|1\rangle$, result is found that the larger dissipations intensities $\gamma_{isol}^{out}$ and $\gamma_{bond}^{out}$, the larger the probability of forming stable hydrogen bond. However, in the second row, when initial state is $|1\rangle|1\rangle|0\rangle$, the opposite result is found. In the first row from panel (a) to panel (d), it is easy to find that the solid dividing line representing the probability of $0.9$ is shifted from top to bottom until it disappears. At the same time, the area of region III of probability with [0.5, 0.9)0.50.9[0.5,\ 0.9)[ 0.5 , 0.9 ) decreases to 00 and the area of region IV of probability with [0.9, 1]0.91[0.9,\ 1][ 0.9 , 1 ] increases to $1$. This means the intensity $\gamma_{phn}^{out}$ promotes the hydrogen bond formation when initial state is $|0\rangle|0\rangle|1\rangle$. Similarly, in the second row from panel (e) to panel (h), the solid dividing line representing the probability of $0.1$ is shifted from right to left. At the same time, the area of region I of probability with [0, 0.1)00.1[0,\ 0.1)[ 0 , 0.1 ) increases and the area of region II of probability with [0.1, 0.5]0.10.5[0.1,\ 0.5][ 0.1 , 0.5 ] decreases. This means intensity $\gamma_{phn}^{out}$ hinders the hydrogen bond formation when initial state is $|1\rangle|1\rangle|0\rangle$. In summary, in the case of $|0\rangle|0\rangle|1\rangle$, the all dissipations intensities have the positive effect on the formation of stable hydrogen bond; in the case of $|1\rangle|1\rangle|0\rangle$, the dissipations intensities have the negative effect on the formation of stable hydrogen bond. Similar to the Fig. 6, when the initial state is $|0\rangle|0\rangle|1\rangle$, the probabilities on the heat maps are all on the range of [0.5, 1]0.51[0.5,\ 1][ 0.5 , 1 ]; When the initial state is $|1\rangle|1\rangle|0\rangle$, the probabilities on the heat maps are all on the range of [0, 0.5)00.5[0,\ 0.5)[ 0 , 0.5 ). This shows that dissipations intensities have a slight effect on evolution, too.
IV.4 The effects of inflows on evolution
Temperature is also an important factor. According to Eq. (7), $\mu_{bond},\ \mu_{isol}$ and $\mu_{phn}$ are described as follows
$\mu_{isol}$ $=\frac{\gamma_{isol}^{in}}{\gamma_{isol}^{out}}=e^{\frac{-\hbar\omega_{isol}}{KT_{isol}}}$ (26a)
$\mu_{bond}$ $=\frac{\gamma_{bond}^{in}}{\gamma_{bond}^{out}}=e^{\frac{-\hbar\omega_{bond}}{KT_{bond}}}$ (26b)
$\mu_{phn}$ $=\frac{\gamma_{phn}^{in}}{\gamma_{phn}^{out}}=e^{\frac{-\hbar\omega_{phn}}{KT_{phn}}}$ (26c)
where $\omega_{isol}=\omega_{bond}=\omega_{phn}$, $\mu_{isol}<1,\ \mu_{bond}<1,\ \mu_{phn}<1$. When $\gamma_{isol}^{out}=\gamma_{bond}^{out}=\gamma_{phn}^{out}=\gamma$, $\mu_{bond},\ \mu_{isol}$ and $\mu_{phn}$ are positively correlated with inflows intensities $\gamma_{isol}^{in}$, $\gamma_{bond}^{in}$ and $\gamma_{phn}^{in}$, respectively; they are also positively correlated with temperature $T_{isol}$, $T_{bond}$ and $T_{phn}$, respectively. For the convenience of research, $\mu_{isol}$, $\mu_{bond}$ and $\mu_{phn}$ can be used to study the effects of inflows or temperatures on evolution. The values of $\mu_{isol}$ and $\mu_{bond}$ range from 00 to $0.9$. The values of $\mu_{phn}$ from the first column to the fourth column correspond to 0, 0.3, 0.600.30.60,\ 0.3,\ 0.60 , 0.3 , 0.6 and $0.9$, respectively. Under different initial states, the dissipative evolution will also eventually reach probabilistic stability after long-term evolution.
Figure 9: (online color) Effect of external impulses of phonons on the hydrogen bond formation. $g_{dist}=g_{prot}=g$, $\gamma_{isol}^{out}=\gamma_{bond}^{out}=\gamma_{phn}^{out}=\gamma$, $\mu_{isol}=\mu_{bond}=0$, $\mu_{phn}=0.01$. The number of phonons varies in different panels: 10 in panel (a), 20 in (b), 30 in (c).
According to the Fig. 8, result is found that no matter the initial state is $|0\rangle|0\rangle|1\rangle$ or $|1\rangle|1\rangle|0\rangle$, the larger $\mu_{isol}$ and the smaller $\mu_{bond}$, the larger the probability of stable hydrogen bond formation. This indicates that whereas $\mu_{bond}$ inhibits the creation of hydrogen bonds, $\mu_{isol}$ encourages them. In the first row from panel (a) to panel (d), it is easy to find that all three solid dividing lines representing the probability of $0.1$, $0.5$ and $0.9$ respectively are shifted from left to right. At the same time, the area of region I of probability with [0, 0.1)00.1[0,\ 0.1)[ 0 , 0.1 ) and the area of region II of probability with [0.1, 0.5)0.10.5[0.1,\ 0.5)[ 0.1 , 0.5 ) both increases, the area of region III of probability with [0.5, 0.9]0.50.9[0.5,\ 0.9][ 0.5 , 0.9 ] and the area of region IV of probability with [0.9, 1]0.91[0.9,\ 1][ 0.9 , 1 ] both decreases. This means the $\mu_{phn}$ hinders the hydrogen bond formation no matter the initial state is $|0\rangle|0\rangle|1\rangle$ or $|1\rangle|1\rangle|0\rangle$. In summary, the $\mu_{isol}$ has the positive effect on the formation of stable hydrogen bond; however, the $\mu_{bond}$ and $\mu_{phn}$ have the negative effect on it. Different to the Figs. 6 and 7, the probabilities on the heat maps are all on the range of [0, 1]01[0,\ 1][ 0 , 1 ]. This shows that inflows intensities (temperature) have a significant effect on evolution than that of other factors (interactions and dissipations).
IV.5 The effects of external impulses on evolution
Figure 10: (online color) Effect of number of phonons on the hydrogen bond formation. $g_{dist}=g_{prot}=g$, $\gamma_{isol}^{out}=\gamma_{bond}^{out}=\gamma_{ph}^{out}=\gamma$, $\mu_{isol}=\mu_{bond}=0$, $\mu_{phn}=0.01$. The number of phonons is within [1, 30]130[1,\ 30][ 1 , 30 ].
@@ IV.5 The effects of external impulses on evolution
As mentioned before, the micro-oscillation of protons will affect the stability of hydrogen bonds. Therefore, in this section, the phonon energy to enhance the effect of this micro-oscillation is strengthened and the changes in the probability of stable hydrogen bond formation is observed. In order to incorporate the effect of external impulses on proton tunneling into the model, the initial number of phonons is changed (in order to reduce interference factors, the inflows intensities $\gamma_{isol}^{in}$ and $\gamma_{bond}^{in}$ is not considered at this time).
[…]
Now consideration is given to the effect of both the initial phonon number and the phonon inflow intensity on the evolution. In Fig. 11, the effect of external impulse and phononic temperature (replaced by $\mu_{phn}$) on the probability of stable hydrogen bond formation is obtained. From the horizontal axis, as the external impulse (the number of phonons in the initial state) increases, the hydrogen bond stability decreases. From the vertical axis, the same conclusion applies to the phononic inflow intensity $\mu_{phn}$. The increase of temperature will make the micro-oscillations of proton more intense, thereby reducing the stability of hydrogen bonds. From the four panels of the first row in Fig. 11, as $\gamma_{bond}^{out}$ increases, the probability of the points in the heat map increases, indicating that $\gamma_{bond}^{out}$ promotes the formation of stable hydrogen bond. In addition, the larger $\gamma_{bond}^{out}$ is, the less obvious the negative effect of $\mu_{phn}$ on stable hydrogen bond formation is. From the four panels of the second row in Fig. 11, as $\gamma_{isol}^{out}$ increases, the probability of the points in the heat map decreases, indicating that $\gamma_{bond}^{out}$ hinders the formation of stable hydrogen bond. In panel (e), $\gamma_{isol}^{out}$ is small enough, the the probability of the points in the heat map hardly changes with the increase of the number of phonons; in panels (f)$\sim$(h), the values of $\gamma_{isol}^{out}$ are large, the probability of the points in each heat map changes with the increase of the number of phonons.

### v3

@@ IV.2 The effect of interactions on evolution
The result of Fig. 6 is shown as heat maps, where the colored dots represent the probability of the stable hydrogen bond formation at $t\to\infty$. Here, $\infty$ is obtained when the system tends to stabilize and the probability of the states approach constant values. In the next subsections, the results of Figs. 7, 8, and 11 are also represented as heat maps at $t\to\infty$.
@@
The effect of interaction strengths $g_{dist}$ and $g_{prot}$ on the evolution and the hydrogen bond formation is obtained in Fig. 6, where the values of $g_{dist}$ and $g_{prot}$ range from $0.5g$ to $2g$. All dissipation intensities are fixed: $\gamma_{isol}^{out}=\gamma_{bond}^{out}=\gamma_{phn}^{out}=\gamma$. Under different initial states, the dissipative evolution will eventually reach probabilistic stability after long-term evolution.
@@
In the first row of Fig. 6, when the initial state is $|0\rangle|0\rangle|1\rangle$, the result demonstrates that the larger the interaction strengths $g_{dist}$ and $g_{prot}$, the lower the probability of the stable hydrogen bond formation. However, in the second row, when the initial state is $|1\rangle|1\rangle|0\rangle$, the opposite result is found. In summary, in the case of $|0\rangle|0\rangle|1\rangle$, the interaction strengths have a negative effect on the formation of the stable hydrogen bond — hindering the hydrogen bond formation; in the case of $|1\rangle|1\rangle|0\rangle$, the interaction strengths have a positive effect on the formation of the stable hydrogen bond — promoting the hydrogen bond formation. Besides, when the initial state is $|0\rangle|0\rangle|1\rangle$, the probabilities on the heat maps are all in the range of [0.5, 1][0.5,\ 1]; when the initial state is $|1\rangle|1\rangle|0\rangle$, the probabilities on the heat maps are all in the range of [0, 0.5)[0,\ 0.5). This is because, according to the results in Fig. 5 (a), when the initial state is $|0\rangle|0\rangle|1\rangle$, the probability of the $|\rm{O}\cdots\rm{H}\rangle$ is greater than $0.8$, and in the first row of Fig. 6 the interaction strengths have a slight effect on the evolution, so the probabilities represented by the points on the first row of the heat maps are almost all around $0.8$. Similarly, according to Fig. 5 (b), when the initial state is $|1\rangle|1\rangle|0\rangle$, the probability of the $|\rm{O}\cdots\rm{H}\rangle$ is less than $0.2$, and in the second row of Fig. 6 the interaction strengths also have a slight effect on the evolution, so the probabilities represented by the points on the second row of the heat maps are almost all around $0.2$.
@@
In this section, the effects of the inflow intensities $\gamma_{phn}^{in}$, $\gamma_{bond}^{in}$, and $\gamma_{isol}^{in}$ are investigated. The first column corresponds to the case that inflows are prohibited: $\mu_{isol}=\mu_{bond}=\mu_{phn}=0$, that is to say, $\gamma_{isol}^{in}=\gamma_{bond}^{in}=\gamma_{phn}^{in}=0$. The second column corresponds to the case that only phonon inflow is permitted: $\mu_{isol}=\mu_{bond}=0,\ \mu_{phn}=0.01$. The third column corresponds to the case that all inflows are permitted: $\mu_{isol}=\mu_{bond}=\mu_{phn}=0.01$. Comparing the second and the third columns with the first column, the result demonstrates that the addition of weak inflows actually promotes the formation of the hydrogen bond, regardless of whether the initial state is $|0\rangle|0\rangle|1\rangle$ or $|1\rangle|1\rangle|0\rangle$. For easy observation, we use the dividing lines. In the first row, from panel (a) to panel (c), it is easy to find that the solid dividing line representing the probability of $0.9$ is shifted to the upper right. At the same time, the area of region III of probability with [0.5, 0.9)[0.5,\ 0.9) decreases and the area of region IV of probability with [0.9, 1][0.9,\ 1] increases. This means the weak inflows promote the hydrogen bond formation when the initial state is $|0\rangle|0\rangle|1\rangle$. Similarly, in the second row from panel (d) to panel (f), the solid dividing line representing the probability of $0.1$ is shifted to the lower left. At the same time, the area of region I of probability with [0, 0.1)[0,\ 0.1) decreases, and the area of region II of probability with [0.1, 0.5][0.1,\ 0.5] increases. This means the weak inflows also promote the hydrogen bond formation when the initial state is $|1\rangle|1\rangle|0\rangle$.
@@
In addition to the above results, the effect of different interaction strengths on the time required for the system to reach stability is also obtained: the smaller the interaction strengths, the longer it takes to reach stability.
IV.3 The effects of dissipation on evolution
Figure 7: (online color) Effect of different dissipation intensities $\gamma_{isol}^{out}$ and $\gamma_{bond}^{out}$, and $\gamma_{phn}^{out}$ on the hydrogen bond formation. $g_{prot}=g_{dist}=g$. The first row of panels is the result of the initial state $|0\rangle|0\rangle|1\rangle$. The second row of panels is the result of the initial state $|1\rangle|1\rangle|0\rangle$. The values of $\gamma_{isol}^{out}$ and $\gamma_{bond}^{out}$ range from $0.5\gamma$ to $2\gamma$, and the values of $\gamma_{phn}^{out}$ from the first column to the fourth column correspond to $\frac{1}{2}\gamma,\ \gamma,\ \frac{3}{2}\gamma$, and $2\gamma$, respectively. The definitions of regions II, III, and IV are all defined in Fig. 6.
The effect of dissipation on evolution is studied. The parameters being studied are $\gamma_{bond}^{out},\ \gamma_{phn}^{out},\ \gamma_{isol}^{out}$. All interaction strengths are fixed: $g_{prot}=g_{dist}=g$. The values of $\gamma_{isol}^{out}$ and $\gamma_{bond}^{out}$ range from $0.5\gamma$ to $2\gamma$. The values of $\gamma_{phn}^{out}$ from the first column to the fourth column correspond to $0.5\gamma,\ \gamma,\ 1.5\gamma$, and $2\gamma$, respectively. Under different initial states, the dissipative evolution will also eventually reach probabilistic stability after long-term evolution.
In the first row of Fig. 7, when the initial state is $|0\rangle|0\rangle|1\rangle$, the result demonstrates that the larger the dissipation intensities $\gamma_{isol}^{out}$ and $\gamma_{bond}^{out}$, the higher the probability of forming the stable hydrogen bond. However, in the second row, when the initial state is $|1\rangle|1\rangle|0\rangle$, the opposite result is found. In the first row, from panel (a) to panel (d), it is easy to find that the solid dividing line representing the probability of $0.9$ is shifted from top to bottom until it disappears. At the same time, the area of region III of probability with [0.5, 0.9)[0.5,\ 0.9) decreases to 0, and the area of region IV of probability with [0.9, 1][0.9,\ 1] increases to $1$. This means the intensity $\gamma_{phn}^{out}$ promotes the hydrogen bond formation when the initial state is $|0\rangle|0\rangle|1\rangle$. Similarly, in the second row from panel (e) to panel (h), the solid dividing line representing the probability of $0.1$ is shifted from right to left. At the same time, the area of region I of probability with [0, 0.1)[0,\ 0.1) increases, and the area of region II of probability with [0.1, 0.5][0.1,\ 0.5] decreases. This means intensity $\gamma_{phn}^{out}$ hinders the hydrogen bond formation when the initial state is $|1\rangle|1\rangle|0\rangle$. In summary, in the case of $|0\rangle|0\rangle|1\rangle$, all dissipation intensities have a positive effect on the formation of the stable hydrogen bond; in the case of $|1\rangle|1\rangle|0\rangle$, all dissipation intensities have a negative effect on the formation of the stable hydrogen bond. Similar to Fig. 6, when the initial state is $|0\rangle|0\rangle|1\rangle$, the probabilities on the heat maps are all in the range of [0.5, 1][0.5,\ 1]; when the initial state is $|1\rangle|1\rangle|0\rangle$, the probabilities on the heat maps are all in the range of [0, 0.5)[0,\ 0.5). This shows that the dissipation intensities have a slight effect on evolution, too.
IV.4 The effects of inflows on evolution
Figure 8: (online color) Effect of different inflow intensities $\mu_{isol}$, $\mu_{bond}$, and $\mu_{phn}$ on the hydrogen bond formation. $g_{prot}=g_{dist}=g$, $\gamma_{isol}^{out}=\gamma_{bond}^{out}=\gamma_{phn}^{out}=\gamma$. The first row of panels is the result of the initial state $|0\rangle|0\rangle|1\rangle$. The second row of panels is the result of the initial state $|1\rangle|1\rangle|0\rangle$. The values of $\mu_{isol}$ and $\mu_{bond}$ range from 0 to $0.9$, and the values of $\mu_{phn}$ from the first column to the fourth column correspond to 0, 0.3, 0.60,\ 0.3,\ 0.6, and $0.9$, respectively. The definitions of regions I, II, III, and IV are all shown in Fig. 6.
Temperature is also an important factor. According to Eq. (7), $\mu_{bond},\ \mu_{isol}$, and $\mu_{phn}$ are described as follows
$\mu_{isol}=\frac{\gamma_{isol}^{in}}{\gamma_{isol}^{out}}=e^{\frac{-\hbar\omega_{isol}}{KT_{isol}}},\ \mu_{bond}=\frac{\gamma_{bond}^{in}}{\gamma_{bond}^{out}}=e^{\frac{-\hbar\omega_{bond}}{KT_{bond}}},\ \mu_{phn}=\frac{\gamma_{phn}^{in}}{\gamma_{phn}^{out}}=e^{\frac{-\hbar\omega_{phn}}{KT_{phn}}},$ (24)
where $\omega_{isol}=\omega_{bond}=\omega_{phn}$, $\mu_{isol}<1,\ \mu_{bond}<1,\ \mu_{phn}<1$. When $\gamma_{isol}^{out}=\gamma_{bond}^{out}=\gamma_{phn}^{out}=\gamma$, $\mu_{bond},\ \mu_{isol}$ and $\mu_{phn}$ are positively correlated with inflow intensities $\gamma_{isol}^{in}$, $\gamma_{bond}^{in}$ and $\gamma_{phn}^{in}$, respectively; they are also positively correlated with temperature $T_{isol}$, $T_{bond}$ and $T_{phn}$, respectively. For the convenience of research, $\mu_{isol}$, $\mu_{bond}$, and $\mu_{phn}$ can be used to study the effects of inflows or temperatures on evolution. The values of $\mu_{isol}$ and $\mu_{bond}$ range from 0 to $0.9$. The values of $\mu_{phn}$ from the first column to the fourth column correspond to 0, 0.3, 0.60,\ 0.3,\ 0.6, and $0.9$, respectively. Under different initial states, the dissipative evolution will also eventually reach probabilistic stability after long-term evolution.
According to Fig. 8, the result demonstrates that no matter if the initial state is $|0\rangle|0\rangle|1\rangle$ or $|1\rangle|1\rangle|0\rangle$, the larger $\mu_{isol}$ and the smaller $\mu_{bond}$, the higher the probability of the stable hydrogen bond formation. This indicates that whereas $\mu_{bond}$ inhibits the creation of hydrogen bonds, $\mu_{isol}$ encourages them. In the first row, from panel (a) to panel (d), it is easy to find that all three solid dividing lines representing the probabilities of $0.1$, $0.5$, and $0.9$, respectively, are shifted from left to right. Simultaneously, the areas under the probability curve corresponding to regions I [0, 0.1)[0,\ 0.1) and II [0.1, 0.5)[0.1,\ 0.5) increase, whereas those for regions III [0.5, 0.9)[0.5,\ 0.9) and IV [0.9, 1][0.9,\ 1] decrease. This means the $\mu_{phn}$ hinders the hydrogen bond formation no matter the initial state is $|0\rangle|0\rangle|1\rangle$ or $|1\rangle|1\rangle|0\rangle$. In summary, the $\mu_{isol}$ has a positive effect on the formation of the stable hydrogen bond; however, the $\mu_{bond}$ and $\mu_{phn}$ have a negative effect on it. Different from Figs. 6 and 7, the probabilities on the heat maps are all in the range of [0, 1][0,\ 1]. This shows that inflow intensities (temperature) have a more significant effect on evolution than that of other factors (interactions and dissipation).
IV.5 The effects of external impulses on evolution
Figure 9: (online color) Effect of external impulses of phonons on the hydrogen bond formation. $g_{dist}=g_{prot}=g$, $\gamma_{isol}^{out}=\gamma_{bond}^{out}=\gamma_{phn}^{out}=\gamma$, $\mu_{isol}=\mu_{bond}=0$, $\mu_{phn}=0.01$. The number of phonons varies in different panels: 10 in panel (a), 20 in (b), and 30 in (c).
Figure 10: (online color) Effect of number of phonons on the hydrogen bond formation. $g_{dist}=g_{prot}=g$, $\gamma_{isol}^{out}=\gamma_{bond}^{out}=\gamma_{ph}^{out}=\gamma$, $\mu_{isol}=\mu_{bond}=0$, $\mu_{phn}=0.01$. The number of phonons is within [1, 30][1,\ 30].
As mentioned before, the micro-oscillation of protons will affect the stability of hydrogen bonds. Therefore, in this section, the phonon energy to enhance the effect of this micro-oscillation is strengthened, and the changes in the probability of the stable hydrogen bond formation are observed. In order to incorporate the effect of external impulses on proton tunneling into the model, the initial number of phonons is changed (in order to reduce interference factors, the inflow intensities $\gamma_{isol}^{in}$ and $\gamma_{bond}^{in}$ are not considered at this time).
@@ IV.5 The effects of external impulses on evolution
Figure 11: (online color) Effect of the initial phonon number and phonon inflow intensity on the hydrogen bond formation. $g_{dist}=g_{prot}=g$, $\mu_{isol}=\mu_{bond}=0$, and the value of $\mu_{phn}$ ranges from 0 to 0.9. The initial state is $|0\rangle|0\rangle|N_{phn}\rangle$, $N_{phn}\in[1,\ 20]$. In the first row, $\gamma_{isol}^{out}=\gamma_{phn}^{out}=\gamma$ is fixed, while $\gamma_{bond}^{out}$ is varied in the four panels: $\gamma_{bond}^{out}=0.1\gamma$ in panel (a), $\gamma_{bond}^{out}=0.4\gamma$ in panel (b), $\gamma_{bond}^{out}=0.7\gamma$ in panel (c), and $\gamma_{bond}^{out}=\gamma$ in panel (d). In the second row, $\gamma_{bond}^{out}=\gamma_{phn}^{out}=\gamma$ is fixed, while $\gamma_{isol}^{out}$ is varied in the four panels: $\gamma_{isol}^{out}=0.1\gamma$ in panel (e), $\gamma_{isol}^{out}=0.4\gamma$ in panel (f), $\gamma_{isol}^{out}=0.7\gamma$ in panel (g), and $\gamma_{isol}^{out}=\gamma$ in panel (h).